OpenStudy (anonymous):

what is the equation of a line parallel to the line whose equation is y=x+2 ?

OpenStudy (amistre64):

parallel lines have the exact same slope. the equation of a straight line is: y=mx+b m=the slope. from y=x+2 you need to remember that there is a (1) that stands in front of the "x" that is not written. Since the "m" = 1; we need to write another stragiht line equation with a slope of (1). y= (1)x + 50 = x+50 is parallel to the original line.

OpenStudy (amistre64):

or better yet; the equation of ANY line to the original you posted is: y= x+ C ; where C= any "real" number.